# SReferences

### v1

@@ Abstract
Semantic segmentation models are typically trained on a fixed set of classes, limiting their applicability in open-world scenarios. Class-incremental semantic segmentation aims to update models with emerging new classes while preventing catastrophic forgetting of previously learned ones. However, existing methods impose strict rigidity on old classes, reducing their effectiveness in learning new incremental classes. In this work, we propose Taxonomy-Oriented Poincaré-regularized Incremental-Class Segmentation (TOPICS) that learns feature embeddings in hyperbolic space following explicit taxonomy-tree structures. This supervision provides plasticity for old classes, updating ancestors based on new classes while integrating new classes at fitting positions. Additionally, we maintain implicit class relational constraints on the geometric basis of the Poincaré ball. This ensures that the latent space can continuously adapt to new constraints while maintaining a robust structure to combat catastrophic forgetting. We also establish eight realistic incremental learning protocols for autonomous driving scenarios, where novel classes can originate from known classes or the background. Extensive evaluations of TOPICS on the Cityscapes and Mapillary Vistas 2.0 benchmarks demonstrate that it achieves state-of-the-art performance. We make the code and trained models publicly available at http://topics.cs.uni-freiburg.de.
@@ I Introduction
Furthermore, most methods are tailored to the highly curated PascalVOC dataset which deviates significantly from densely annotated automated driving scenarios. While two to three object categories appear per image in PascalVOC, driving datasets typically contain over twenty different object categories, and fewer pixels are assigned to the background class. Additionally, all CISS methods assume that new classes originate from the prior background. This scenario is unrealistic for automated driving as a change in requirements for navigation could also entail bifurcations of previously observed classes for better decision-making, e.g. a model is initially trained to uniformly segment humans but later this ability needs to be extended to distinguish different vulnerable road users.
@@
In this work, we introduce taxonomy-aware continual semantic segmentation for automated driving scenarios. Our proposed Taxonomy-Oriented Poincaré-regularized Incremental Class Segmentation (TOPICS) approach enforces features conform to taxonomy-tree structures in hyperbolic space. As a result, the overall class distribution is rigid and new classes are appended at fitting positions in Fig. 1. This supervision also provides plasticity for old classes as the positions of ancestors are updated according to new classes. We show that this regularization is beneficial for open-world scenarios since it includes aspects of plasticity and rigidity. To further avoid catastrophic forgetting, we incorporate pseudo-labeling of the background and relation constraints of old class hyperplanes in TOPICS. We argue that semantic classes inherit relations that go beyond the defined class taxonomy such as similar appearances or contexts. Those relations can be observed as close mappings in latent space. We ensure consistency of the relations between prior class hyperplanes to increase the rigidity of our model. Lastly, we constrain features to have equidistant radii to maintain constant scarcity. We ensure that prior class features can only move in a circular direction around the hyperbolic center. Accordingly, new classes cannot result in a latent space shift of the complete taxonomy in their favor. We perform extensive evaluations of TOPICS on the standard Cityscapes [Cordts2016Cityscapes] and Mapillary Vistas 2.0 [neuhold2017mapillary] benchmarks where it sets the new state-of-the-art.
@@ II Related Work
In this section, we summarize existing works in class-incremental semantic segmentation, hyperbolic neural networks, and hierarchical learning.
@@
Class-Incremental Semantic Segmentation: CISS methods rely on data replay, expansion, or distillation to avoid catastrophic forgetting. Prior data is recreated with GANs [maracani2021recall] or a small subset of prior data is stored in memory-buffers [cha2021_ssul]. To further reduce memory constraints, prior work stores selected feature representations instead of raw data [chen2023STAR]. Expansion-based methods dedicate separate network components for particular semantic knowledge. For example, one branch of parallel convolutions adapts to the new data and is merged into the frozen branch after every incremental step [zhang2022representation]. Follow-up work extends this idea by fusing only endpoints of a trainable and frozen model in combination with distillation [xiao2023endpoints]. Distillation approaches maintain prior model weights to restrain the current model for equivalent responses to the input data [zhou2023deep]. The pioneering approach MiB [cermelli2020mib] relates prior background logits to the combination of novel-class and background logits in the new model. This method is enhanced with gradient-based attribution weight initialization which identifies relevant classifier weights for novel classes from prior weights of the background class [goswami2023attribution]. On the other hand, PLOP [Douillard2020PLOPLW] labels the background with prior model predictions and distills pooled intermediate feature representations. Subsequent work focuses on learning an enhanced weighting term for distillation [wang24] or adapting this principle to transformer architectures [cermelli2023comformer, qiu2023sats]. The method SATS [qiu2023sats] also highlights the benefit of relation distillation between self-attention vectors in a SegFormer model. This weaker constraint allows the model to avoid forgetting while not constraining its plasticity. Prior work also trains segmentation models with sigmoid activation and binary cross-entropy loss as the instability of softmax activations hinders incremental learning [cha2021_ssul]. DKD [baek2022_dkd] further combines this approach with decoupled knowledge distillation while other approaches completely freeze the feature extractor and segment unknown background classes with saliency detectors [cha2021_ssul] or pre-trained models [zhang22_microseg].
@@
However, all observed methods focus on novel classes from the background which significantly hinders their applicability in real-world scenarios where incremental learning could also entail a refinement of known classes. In this paper, we propose to simultaneously benchmark CISS methods for incremental learning from known classes and the background in the context of autonomous driving. Further, CISS methods do not utilize semantic relationships between classes to balance plasticity and rigidity. We hypothesize that a hierarchical mapping of class features facilitates learning new classes while it constrains forgetting old classes.
Hyperbolic Neural Networks: Hyperbolic neural networks have been first proposed to capture tree-like structures in text and graphs. For CNNs, the hyperbolic classification is modeled as a prototype-based approach [lang2022hyperbolic, cui2023_hypcil] or multinomial logistic regression as proposed by [ganea2018hnn]. For the first, a cone entailment loss enforces all descendant prototypes to lie in the same geometric cone [desai2023meru]. In multinomial logistic regression, semantic classes are geometrically interpreted as hyperplanes, and hierarchies are explicitly modeled with a hierarchical softmax [atigh2022hyperbolic] or cosine margins [xu2023hier]. While hyperbolic neural networks have been extensively explored in image classification [xu2023hier, Spengler2023PoincarR] and metric learning [desai2023meru], few prior works focus on semantic segmentation. Atigh et al. [atigh2022hyperbolic] first shows the potential of applying hyperbolic multinomial regression for image segmentation. Follow-up work highlights the dense calibration capabilities of this network [weber2024flattening] and leverages it for active learning [franco2024hyperbolic]. One pioneering work also explores hyperbolic spaces for class-incremental image classification and highlights the positive impact of geometric-induced clear decision boundaries [cui2023_hypcil]. Motivated by them, our work is the first to explore hyperbolic spaces for CISS.
@@
Hierarchical Learning: Hierarchical segmentation methods outperform flat classifiers on leaf categories by modeling the semantic hierarchy of classes in features [chen2023taxonomic] or combining logits with those of ancestor classes [liulei24, atigh2022hyperbolic]. While hierarchical semantic segmentation primarily addresses closed class settings, prior work in image classification concentrates on taxonomic class-incremental learning with network expansion [chen2023taxonomic] and replay-buffers [lin2022continual]. Lin et al. first focused on taxonomic incremental semantic segmentation [lin2022continual]. However, they require all ancestor classes to be present in the base training and allow the usage of all history data which contradicts CISS principles. We introduce a more realistic form of taxonomic class-incremental semantic segmentation which allows increments from both background and known classes and complies with other CISS task definitions.
@@ III Technical Approach
Figure 2: During base training of TOPICS, features are mapped onto the Poincaré ball before the class hierarchy is explicitly enforced with $\mathcal{L}_{hier}$. In incremental steps, the old model is used to generate pseudo-labels of old classes and to regularize the last layer’s weights with $\mathcal{L}_{rel}$ and feature radii with $\mathcal{L}_{dist}$.
@@
In this section, we first introduce taxonomic CISS. We then present our TOPICS approach which is tailored to incrementally learn new classes from either background or prior known classes as shown in Fig. 2. We leverage the class taxonomy and implicit relations between prior classes to avoid catastrophic forgetting in incremental learning steps. We first train the model on the base dataset. The class hierarchy is explicitly enforced in the final network layer which is mapped in hyperbolic space. This geometric space ensures that classes are equidistant to each other irrespective of their hierarchy level which facilitates learning tree-like class hierarchy structure. During the incremental steps, we leverage the old model’s weights to create pseudo-labels for the background and employ scarcity and relation regularization losses to maintain important relations of old classes while learning the novel classes in a supervised manner.
@@ III-C Hierarchical Segmentation
which penalizes hierarchical-inconsistent predictions for every class. For a correct prediction $l$, this loss penalizes the smallest logit of its ancestors $\mathcal{A}_{v}$. On the other hand, for wrong predictions $1-l$, $\mathcal{L}^{\mathrm{TM}}$ punishes the maximum logits of its descendant $\mathcal{D}_{v}$ to reduce the score of complete root-to-leaf branches.
@@ III-E Hyperbolic Distance Correlation
14-1 (6 tasks) 10-1 (10 tasks) 7-4 (4 tasks)h 7-18 (2 tasks)h
@@ IV Experimental Evaluation
In this section, we present quantitative and qualitative results of TOPICS on nine CISS settings in addition to a comprehensive ablation study to underline the importance of our contributions. Further, we detail the applied CISS settings and the training protocol that we employ.
@@ IV-A Datasets
We evaluate TOPICS on the Cityscapes [Cordts2016Cityscapes] and Mapillary Vistas 2.0 [neuhold2017mapillary] datasets. For both datasets, we define CISS protocols where incremental classes either primarily originate from the background or known classes. We only consider the more realistic case of overlapped CISS where image pixels can belong to old, current, or future classes at any timestep. The Cityscapes dataset consists of 19 semantic classes in addition to a void class. For CISS from the background, we adapt the 14-1 (6 tasks) and 10-1 (10 tasks) setting as proposed in [goswami2023attribution]. The first 10 or 14 classes are learned during base training while one class is added per incremental step. The task count includes base training as the first task. For CISS from known classes, we learn 7 base classes that correspond to the official sub-categories defined for Cityscapes and increment the model in a 7-4 (4 tasks) or 7-18 (2 tasks) manner.
@@ IV-B Experimental Setup
In line with prior work [Douillard2020PLOPLW, cermelli2020mib, zhang22_microseg], we use the DeepLabV3 model with the ResNet-101 backbone which is pre-trained on ImageNet for all the experiments. We employ the Geoopt library [geoopt2020kochurov] to project the Euclidean features to a Poincaré ball with $c=2.0$ (different curvatures are explored in Sec. LABEL:sec:curv). Further, we follow the Möbius approximation defined in [atigh2022hyperbolic] for more efficient computations. We train TOPICS for 60 epochs per task with batch size $24$ for Cityscapes and $16$ for Mapillary Vistas 2.0 using the Riemannian SGD optimizer with momentum of $0.9$ and weight decay of $0.0001$. We use a poly learning rate scheduler with initial learning rates of $0.05$ for base training and $0.01$ in all incremental steps. We additionally ablate lower learning rates in Sec. IV-D2. For the hierarchical loss function, we set the hyper-parameters to $\alpha=5$ and $\beta=1$ and ablate different hierarchical functions in Sec. LABEL:sec:hh. For Mapillary Vistas 2.0, we rescale the longest size to 2177 pixels before taking a non-empty crop of (1024,1024) and horizontal flipping. On the other hand, we train on random non-empty crops of (512,1024) with horizontal flipping for Cityscapes. Non-empty cropping biases image crops to include labeled masks (i.e. new classes) which could be neglected when applying random cropping.
@@ IV-C Quantitative Results
We compare TOPICS with five state-of-the-art CISS methods: PLOP [Douillard2020PLOPLW], MiB [cermelli2020mib], MiB+AWT [goswami2023attribution], DKD [baek2022_dkd] and MicroSeg [zhang22_microseg]. For each method, we use the respective author’s published code and use the same augmentations outlined in Sec. IV-B. For Cityscapes, we train the method PLOP on $512\times 512$ crops as the method is restricted to squared input images. We evaluate the models using the mean intersection-over-union (mIoU) metric. Specifically, we evaluate the mIoU over all the base classes ($\mathcal{C}_{1}$) and novel classes ($\mathcal{C}_{2:T}$) separately as an indication of rigidity and plasticity. We present the results on Cityscapes in Tab. I. On this dataset, TOPICS outperforms all baselines by at least $9.88$pp on the CISS from the background. While the difference in base IoU measures $4.2$pp, our method significantly exceeds the benchmarks by at least $16.9$pp in terms of novel IoU. This finding emphasizes that a balance between plasticity and rigidity is crucial to achieving superior results for class incremental learning. Further, we note the largest performance difference on incremental scenarios from known classes where TOPICS exceeds the best baseline by $29.94$pp in mIoU. Consequently, we highlight the versatility of our method to balance plasticity and rigidity in all tested CISS settings. DKD [baek2022_dkd] achieves the highest benchmark on CISS from the background in our setting but does not generalize on the bifurcation of previously observed classes which we reason with frozen layers in incremental steps. The MicroSeg approach also does not result in a favorable performance on Cityscapes and Mapillary Vistas 2.0 which is caused by low-quality proposals retrieved from the COCO dataset in addition to fully freezing the backbone. Subsequently, we present the mIoU over different increments for the 10-1(10 tasks) scenario on Cityscapes in Fig. 5. We note that for most benchmarks the mIoU increasingly deteriorates after incremental step 3. On the other hand, TOPICS achieves a constant performance from step 1 to step 6 after which the performance again stabilizes. Therefore, our approach supremely maintains base as well as incremental class knowledge while not restricting the adaptability to new classes.
002222444466668888303030304040404050505050606060607070707080808080Incremental stepmIoU [%]TOPICSMiB[cermelli2020mib]MiB+AWT [goswami2023attribution]PLOP [Douillard2020PLOPLW]DKD [baek2022_dkd]Microseg [zhang22_microseg]
Figure 5: Performance at every increment on Cityscapes 10-1 (10 task) setting.

### v2

@@ Abstract
Semantic segmentation models are typically trained on a fixed set of classes, limiting their applicability in open-world scenarios. Class-incremental semantic segmentation aims to update models with emerging new classes while preventing catastrophic forgetting of previously learned ones. However, existing methods impose strict rigidity on old classes, reducing their effectiveness in learning new incremental classes. In this work, we propose Taxonomy-Oriented Poincaré-regularized Incremental-Class Segmentation (TOPICS) that learns feature embeddings in hyperbolic space following explicit taxonomy-tree structures. This supervision provides plasticity for old classes, updating ancestors based on new classes while integrating new classes at fitting positions. Additionally, we maintain implicit class relational constraints on the geometric basis of the Poincaré ball. This ensures that the latent space can continuously adapt to new constraints while maintaining a robust structure to combat catastrophic forgetting. We also establish eight realistic incremental learning protocols for autonomous driving scenarios, where novel classes can originate from known classes or the background. Extensive evaluations of TOPICS on the Cityscapes and Mapillary Vistas 2.0 benchmarks demonstrate that it achieves state-of-the-art performance. We make the code and trained models publicly available at http://topics.cs.uni-freiburg.de [^1] Code will be published upon acceptance. The authors thank Kshitij Sirohi for technical discussions..
@@ I Introduction
Furthermore, most methods are tailored to the highly curated PascalVOC dataset which deviates significantly from densely annotated automated driving scenarios. While two to three object categories appear per image in PascalVOC, driving datasets typically contain over twenty different object categories, and fewer pixels are assigned to the background class. Additionally, all CISS methods assume that new classes originate from the prior background. This scenario is unrealistic for automated driving as a change in requirements for navigation could also entail bifurcations of previously observed classes for better decision-making, e.g., a model is initially trained to uniformly segment humans but later this ability needs to be extended to distinguish different vulnerable road users.
@@
In this work, we introduce taxonomy-aware continual semantic segmentation for automated driving scenarios. Our proposed Taxonomy-Oriented Poincaré-regularized Incremental Class Segmentation (TOPICS) approach enforces features conform to taxonomy-tree structures in hyperbolic space. As a result, the overall class distribution is rigid and new classes are appended at fitting positions in Fig. 1. This supervision also allows old classes to evolve as the positions of ancestors are updated according to new classes which is a unique property of our method. We find that applying our distinct taxonomy-aware structure is beneficial for open-world scenarios since it effortlessly unites plasticity and rigidity. To further avoid catastrophic forgetting, we incorporate pseudo-labeling of the background and relation constraints of old class hyperplanes in TOPICS. We argue that semantic classes inherit relations that go beyond the defined class taxonomy such as similar appearances or contexts. We preserve consistency of these relations by maintaining relative mappings of prior class hyperplanes. Lastly, we constrain features to have equidistant radii to maintain constant scarcity. We ensure that prior class features can only move in a circular direction around the hyperbolic center. Accordingly, new classes cannot result in a latent space shift of the complete taxonomy in their favor. We perform extensive evaluations of TOPICS on the standard Cityscapes [Cordts2016Cityscapes] and Mapillary Vistas 2.0 [neuhold2017mapillary] benchmarks where it sets the new state-of-the-art.
@@ II Related Work
In this section, we summarize existing works in class-incremental semantic segmentation, hierarchical learning, and hyperbolic neural networks.
@@
Class-Incremental Semantic Segmentation (CISS) methods rely on data replay, expansion, or distillation to avoid catastrophic forgetting. In data replay methods, a subset of raw data [cha2021_ssul] or selected feature representations are stored in memory-buffers [chen2023STAR]. Further, replay data can also be recreated with GANs [maracani2021recall]. Expansion-based methods dedicate separate network components for particular semantic knowledge in the form of parallel convolutions [zhang2022representation] or network branches [xiao2023endpoints]. The trainable branch is merged into the frozen branch after every incremental step to maintain consistent network sizes  [zhang2022representation, xiao2023endpoints]. Distillation approaches maintain prior model weights to restrain the current model for equivalent responses to the input data [zhou2023deep]. The pioneering approach MiB [cermelli2020mib] relates prior background logits to the combination of novel-class and background logits in the new model. This method is enhanced with gradient-based attribution weight initialization which identifies relevant classifier weights for novel classes from prior weights of the background class [goswami2023attribution]. On the other hand, PLOP [Douillard2020PLOPLW] labels the background with prior model predictions and distills pooled intermediate feature representations. Subsequent work focuses on learning an enhanced weighting term for distillation [wang24] or adapting this principle to transformer architectures [cermelli2023comformer, qiu2023sats]. The method SATS [qiu2023sats] also highlights the benefit of relation distillation between self-attention vectors in a SegFormer model. This weaker constraint allows the model to avoid forgetting while not constraining its plasticity. Prior work also trains segmentation models with sigmoid activation and binary cross-entropy loss as the instability of softmax activations hinders incremental learning [cha2021_ssul]. DKD [baek2022_dkd] further combines this approach with decoupled knowledge distillation while other approaches completely freeze the feature extractor and segment unknown background classes with saliency detectors [cha2021_ssul] or pre-trained models [zhang22_microseg].
@@
However, all observed methods focus on novel classes from the background which significantly hinders their applicability in real-world scenarios where incremental learning could also entail a refinement of known classes. In this paper, we propose to simultaneously benchmark CISS methods for incremental learning from known classes and the background in the context of autonomous driving. Further, CISS methods do not include any semantic relationships between classes to balance plasticity and rigidity. We hypothesize that a hierarchical mapping of class features facilitates learning new classes while it constrains forgetting old classes.
@@
Hierarchical Learning methods outperform flat classifiers by modeling the semantic hierarchy in features [chen2023taxonomic] or combining logits with those of ancestor classes [liulei24, atigh2022hyperbolic]. While hierarchical semantic segmentation primarily addresses closed-set cases, prior work in image classification concentrates on taxonomic class-incremental learning with network expansion [chen2023taxonomic] and replay-data [lin2022continual]. Lin et al. first focused on taxonomic incremental semantic segmentation [lin2022continual]. However, they require all ancestor classes to be present in base training and allow using all history data which contradicts CISS principles. We introduce a more realistic form of taxonomic class-incremental semantic segmentation which allows increments from both background and known classes and complies with other CISS task definitions.
Hyperbolic Neural Networks optimally capture tree-like structures in text and graphs [ganea2018hnn]. For CNNs, the hyperbolic classification is modeled as a prototype-based approach [lang2022hyperbolic, cui2023_hypcil] or multinomial logistic regression [ganea2018hnn]. For the first, a cone entailment loss enforces all descendant prototypes to lie in the same geometric cone [desai2023meru]. In multinomial logistic regression, semantic classes are geometrically interpreted as hyperplanes, and hierarchies are explicitly modeled with a hierarchical softmax [atigh2022hyperbolic] or cosine margins [xu2023hier]. While hyperbolic neural networks have been extensively explored in image classification [xu2023hier, Spengler2023PoincarR] and metric learning [desai2023meru], few prior works focus on semantic segmentation. Atigh et al. [atigh2022hyperbolic] first shows the potential of hyperbolic multinomial regression for image segmentation. Follow-up work highlights the dense calibration capabilities of this network [weber2024flattening]. One pioneering work  [cui2023_hypcil] incorporates hyperbolic metric learning to enhance Euclidean-based knowledge retention for incremental image classification, which highlights the need for unique knowledge preservation methods that fully utilize the hyperbolic space. Motivated by them, our work is also the first to explore hyperbolic spaces for CISS.
@@ III Technical Approach
Figure 2: During base training of TOPICS, features are mapped onto the Poincaré ball before the class hierarchy is explicitly enforced with $\mathcal{L}_{hier}$. In incremental steps, the old model is used to generate pseudo-labels of old classes (PL) and to regularize the last layer’s weights with $\mathcal{L}_{rel}$ and feature radii with $\mathcal{L}_{dist}$.
@@
In this section, we first introduce taxonomic CISS. We then present our TOPICS approach which is tailored to incrementally learn new classes from either background or prior known classes as shown in Fig. 2. We leverage the class taxonomy and implicit relations between prior classes to avoid catastrophic forgetting in incremental learning steps. We first train the model on the base dataset. The class hierarchy is explicitly enforced in the final network layer which is mapped in hyperbolic space. This geometric space ensures that classes are equidistant to each other irrespective of their hierarchy level which facilitates learning tree-like hierarchy structure. During the incremental steps, we leverage the old model’s weights to create pseudo-labels for the background and employ the class taxonomy, scarcity, and relation regularization losses to maintain explicit and implicit relations of old classes while learning the novel classes.
@@ III-C Hierarchical Segmentation
which penalizes hierarchical-inconsistent predictions for every class. For a correct prediction $l$, this loss penalizes the smallest logit of its ancestors $\mathcal{A}_{v}$. Conversely, for wrong predictions $1-l$, $\mathcal{L}^{\mathrm{TM}}$ punishes the maximum logits of its descendant $\mathcal{D}_{v}$ to reduce the score of complete root-to-leaf branches.
@@ III-E Hyperbolic Distance Correlation
14-1 (6 tasks) 10-1 (10 tasks) 7-5 (3 t.) -4 (2 t.)h 7-18 (2 tasks)h
@@ IV Experimental Evaluation
In this section, we present quantitative and qualitative results of TOPICS on eight CISS settings in addition to a comprehensive ablation study to underline the importance of our contributions. Further, we detail the applied CISS settings and the training protocol that we employ.
@@ IV-A Datasets
We evaluate TOPICS on the Cityscapes [Cordts2016Cityscapes] and Mapillary Vistas 2.0 [neuhold2017mapillary] datasets. For both datasets, we define CISS protocols where incremental classes either primarily originate from the background or known classes. We only consider the more realistic case of overlapped CISS, where image pixels can belong to old, current, or future classes at any timestep. The Cityscapes dataset consists of 19 semantic classes in addition to a void class. For CISS from the background, we adapt the 14-1 (6 tasks) and 10-1 (10 tasks) setting as proposed in [goswami2023attribution]. The first 14 or 10 classes ($\mathcal{C}^{1}$) are learned during base training, while in this case, one class ($\mathcal{C}^{T}$) is added per incremental step. The task count ($T$) includes base training as the first task. For CISS from known classes, we learn 7 base classes that correspond to the official sub-categories defined for Cityscapes and increment the model in a 7-18 (2 tasks) or 7-5 (3 tasks) -4 (2 tasks) manner.
@@ IV-B Experimental Setup
In line with prior work [Douillard2020PLOPLW, cermelli2020mib, zhang22_microseg], we use the DeepLabV3 model with the ResNet-101 backbone which is pre-trained on ImageNet for all the experiments. We employ the Geoopt library [geoopt2020kochurov] to project the Euclidean features to a Poincaré ball with $c=2.0$ (different curvatures are explored in Sec. LABEL:sec:curv). Further, we follow the Möbius approximation defined in [atigh2022hyperbolic] for more efficient computations. We train TOPICS for 60 epochs per task with batch size $24$ for Cityscapes and $16$ for Mapillary Vistas 2.0 using the Riemannian SGD optimizer with momentum of $0.9$ and weight decay of $0.0001$. We use a poly learning rate scheduler with initial learning rates of $0.05$ for base training and $0.01$ in all incremental steps. We additionally ablate lower learning rates in Sec. IV-D2. For the hierarchical loss function, we set the hyper-parameters to $\alpha=5$ and $\beta=1$ and ablate different hierarchical functions in Sec. LABEL:sec:hh. For Mapillary Vistas 2.0, we rescale the longest size to 2177 pixels before taking a non-empty crop of (1024,1024) and horizontal flipping. On the other hand, we train on random non-empty crops of (512,1024) with horizontal flipping for Cityscapes. Non-empty cropping biases image crops to include labeled masks (i.e., new classes) which could be neglected when applying random cropping.
@@ IV-C Quantitative Results
We compare TOPICS with five state-of-the-art CISS methods: PLOP [Douillard2020PLOPLW], MiB [cermelli2020mib], MiB+AWT [goswami2023attribution], DKD [baek2022_dkd] and MicroSeg [zhang22_microseg]. For each method, we use the respective author’s published code (including hyperparameters) and use the same augmentations outlined in Sec. IV-B. For Cityscapes, we train the PLOP [Douillard2020PLOPLW] on $512\times 512$ crops as the method is limited to squared input images. We evaluate the models using the mean intersection-over-union (mIoU) metric. Specifically, we evaluate the mIoU over all the base classes ($\mathcal{C}_{1}$) and novel classes ($\mathcal{C}_{2:T}$) separately as an indication of rigidity and plasticity. We present the results on Cityscapes in Tab. I. On this dataset, TOPICS outperforms all baselines by at least $9.88$pp on the CISS from the background. While the difference in base IoU measures $4.2$pp, our method significantly exceeds the benchmarks by at least $16.9$pp in terms of novel IoU. This finding emphasizes that a balance between plasticity and rigidity is crucial to achieving superior results for class incremental learning. Further, we note the largest performance difference on incremental scenarios from known classes where TOPICS exceeds the best baseline by $29.94$pp in mIoU. Thus, we highlight the versatility of our method to balance plasticity and rigidity in all tested CISS settings. DKD [baek2022_dkd] achieves the highest benchmark on CISS from the background in our setting but does not generalize on the bifurcation of previously observed classes which we reason with frozen layers in incremental steps. MicroSeg [zhang22_microseg] also does not result in a favorable performance on Cityscapes and Mapillary Vistas 2.0 which is caused by low-quality proposals retrieved from the COCO dataset and fully freezing the backbone. Further, we present the mIoU over different increments for the 10-1 (10 tasks) scenario on Cityscapes in Fig. 5. We note that for most benchmarks, the mIoU increasingly deteriorates after increment $T^{3}$. Conversely, TOPICS achieves a constant performance from $T^{1}$ to $T^{6}$, after which the performance again stabilizes. Therefore, our approach supremely maintains prior class knowledge without restricting adaptability to new classes.
1111333355557777999910101010303030304040404050505050606060607070707080808080Incremental stepmIoU [%]TOPICSTOPICSb2MiB[cermelli2020mib]MiB+AWT [goswami2023attribution]PLOP [Douillard2020PLOPLW]DKD [baek2022_dkd]Microseg [zhang22_microseg]
Figure 5: Performance at every increment on Cityscapes 10-1 (10 task) setting. TOPICSb2 is trained for half iterations in base training to illustrate knowledge retention with lower base training performance.